Find the general solution of the given differential equation

(x+!) dy/dx + (x+2)y = 2xe^-x

y = ______

Determine whether there are any transient terms in the general solution.
1
Expert's answer
2020-07-21T13:42:02-0400


Divide throughout by "(x+1)" :

"\frac{dy}{dx} + \frac{(x+2)}{(x+1)} y= \frac{2xe^{-x }}{(x + 1)}"


"IF = e^{\int \frac{(x+2)}{(x+1)}dx}"


"IF = e^{\int 1+ \frac{1}{(x+1)}dx}"


"IF = e^{x+ln|x+1|}"


"IF = (x + 1)e^{x}"

Multiply throughout by the Integrating factor :


"(x + 1)e^{x} * [\frac{dy}{dx} + \frac{(x+2)}{(x+1)} y = \frac{2xe^{-x} }{ (x + 1)}]"


"\frac{d}{dx}(y (x + 1)e^{x}) = (x + 1)e^{x} \frac{2xe^{-x}}{(x + 1)}"


"\frac{d}{dx}(y (x + 1)e^{x}) =2x"


Integrating :


"y(x + 1)e^{x} = x^2 + C"


"y = \frac{x^2 + C}{(x+1)e^x}"


Clearly, since we have the "x + 1" in denominator, x cannot be equal to -1

This splits number line into intervals "(- \infin , -1)" and "(-1 , \infin)"

out of which "(-1 , \infin)" is LARGER

So, "(-1 , \infin)" "\implies" ANSWER


Both terms, "\frac{x^2}{(x+1)e^x}" and "\frac{C}{(x+1)e^x}" are transient because as "x\to \infin" , both of these terms tend to 0.


Answer: "\frac{x^2}{(x+1)e^x},\frac{C}{(x+1)e^x}" .